Name the values of the characteristics of the following quadratic equations
2. f(x) = 3(x – 5)^2 + 2 3. g(x) = -4x^2 – 8x – 3
axis of symmetry,
vertex
Max or Min point
Range
x-intercepts
amplitude

Answer by stanbon(75887)

2. f(x) = 3(x – 5)^2 + 2
axis of symmetry: x=5
vertex: (5,2)
Max or Min point: Min at the vertex
Range: y>=2
x-intercepts:
Let y=0
Then (x-5)^2=-2/3
Since it is negative there are no x-intercepts
amplitude: N/A
----------------
3. g(x) = -4x^2 – 8x – 3
g(x)+3-4 = -4(x^2+2x+1)
g(x)-1 = -4(x+1)^2
axis of symmetry:x=-1
vertex:(-1,1)
Max or Min point:Max at the vertex
Range : Y<=1
x-intercepts:
Let y=0
(x+1)^2=1/4
(x+1)=+1/2 or (x+1)=-1/2
x=-1/2 or x=-3/2
amplitude: N/A
-------------
Cheers,
Stan H.